Circular No. 8743

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2006 Q2 (LONEOS) An apparently asteroidal object discovered in the course of the LO-

NEOS survey (discovery observation tabulated below), and posted on the ‘NEO Confirmation Page’, has been found to show cometary appearance by other CCD astrometrists. P. Birtwhistle (Great Shefford, Berkshire, U.K., 0.40-m f/6 Schmidt-Cassegrain reflector) reports that his stack of images (total exposure time 44 min) from Aug. 29.9 UT shows that the object has a 9′′-diameter coma (FWHM 4′′.7) and a faint tail ∼ 30′′ long (initially in p.a. 180^{o} but then curving back after 7′′ in p.a. 230^{o}). Several long exposures by J. Young with the 0.61-m f/16 Cassegrain reflector at Table Mountain around Aug. 30.2 show a coma with diameter 8′′ and a broad tail 26′′ long in p.a. 215^{o}. Shorter exposures by Birtwhistle and Young show the comet’s head to be essentially stellar in appearance. Long exposures by Young on Aug. 31.22–31.30 continue to show a tail ∼ 26′′ long in p.a. ∼ 215^{o}, with a coma diameter of 4′′–5′′. Images taken by L. Buzzi (Varese, Italy, 0.60-m f/4.64 reflector) around Aug. 31.0 show a broad, short tail about 10′′ long in p.a. ∼ 180^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Aug. 29.27246  22^{h}38^{m}17.^{s}76  +13^{o}59′30′′.9  18.2

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2006-Q67.

T = 2006 Sept. 2.755 TT  ω = 96.^{o}607
e = 0.58860  Ω = 245.374 }2000.0
q = 1.33560 AU  i = 5.339
a = 3.24643 AU n^{o} = 0.168498 P = 5.85 years

SUPERNOVA 2006eo IN MCG –03-12-15 Further to IAUC 8741, H. Khandrika and W. Li report the LOSS

discovery of an apparent supernova on unfiltered KAIT images taken on Aug. 26.51 (at mag 18.2) and 28.50 UT (mag 18.6). SN 2006eo is located at α = 4^{h}33^{m}21^{s}.94, δ = −18^{o}40′36′′.8 (equinox 2000.0), which is 7′′.1 west and 1′′.3 north of the nucleus of the galaxy MCG –03-12-15. A KAIT image taken on Feb. 9.20 showed nothing at this position (limiting mag 19.0).

2006 August 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green